Three people are moving with velocity v , 2v and 3v respectively. They see rain falling at angles alpha , beta and gamma respectively. What is the relation between alpha beta and gamma ?
1
Expert's answer
2019-08-15T09:25:45-0400

Let the speed of rain will be V.

We have

"\cot{\alpha}=\frac{v}{V}"

"\cot{\beta}=\frac{2v}{V}"

"\cot{\gamma}=\frac{3v}{V}"

the relation between alpha beta and gamma:


"\cot{\alpha}:\cot{\alpha}:\cot{\alpha}=1:2:3"